Firms U and L are in the same risk class and that both have EBIT = $1,000,000. Firm U uses no debt financing and its cost of equity is KsU=15%. Firm L has $2 million of debt outstanding at a cost of Kd = 5%. There are no taxes and MM assumptions hold.

  1. Find V, S, Ks, and WACC for firms U and L.
  2. Using the data given above, but now assuming that firms L and U are both subject to a 40% corporate tax rate, repeat the analysis under the MM with-tax model.
  3. Now suppose investors are subject to the following tax rates: TD=20% and TS=10%. What is the gain from leverage according to the Miller’s Model?
  4. How does this gain compare to the gain in the MM model with corporate taxes?
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy ▼ ,_a. ars-函Merge & Center, $, % , 弼,8 C Paste B l u. Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard SB186 RZ Font Alignment Number Styles Cells SA SB SC SD SE SF SG SH SI SJ SK SM 174 175 176 CASE 1 MMAPPROACH WITH NO TAXES UNLEVERED FIRM VU VALUE OF UNLEVERED FIRM EBIT/ke VU1000000/0.15 VU = 178 179 180 181 182 183 184 185 186 187 188 189 190 191 192 4 H KE CAPM UTILITY, SH 6666666.67 FOR UNLEVERED FIRM V = S= 6,666,666.67 Ks = WACC = 15% LEVERED FIRM WHEN TAXES ARE ABSENT VALUE OF LEVERED FIRM = VALUE OF UNLEVERED FIRM VALUE OF LEVERED FIRM 6666666.67 VL= S-VL-DEBT = 6666666.67-2000000 = 4,666,666.67 66,66,666.67 beta bond c future INDEX INTL CAP BUD ING PV, FV, ANNUITY DIR cleanYIELD bond cap stru wACC | 福 130% RESI ex di㈠ rences: CF135Home nert Page Layout Formulas Data Review View dd-Ins s Cut ta copy. Σ AutoSum ー E ゴWrap Text ,_a. ars-函Merge & Center, $, % , 弼,8 C Paste B 1 u. Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard SB196 RZ Font Alignment Number Styles Cells Edting SA SB SC SD SE SF SG SH SI SJ SK SL SM 193 194 195 196 197 198 199 200 201 202 203 204 205 206 207 208 209 210 211 4 H KE CAPM UTILITY, SH Ks Ko(Ko-Kd) (D/s) Ks-15% + (15%-5%)(2000000/4666666.67) 19.29% (4666666.67/6666666.67)(19.29%) + (2000000/6666666.67)(5%) WACC- WACC- 15.00% CASE 2 MMAPPROACH WITH TAXES VU VALUE OF UNLEVERED FIRM EBIT(1-t)/ke VU= 1000000(1-0.40/0.15 4000000.00 FOR UNLEVERED FIRM V = S= 4000000 Ks-WACC-15% beta bond c future INDEX INTL CAP BUD ING PV, FV, ANNUITY DIR cleanYIELD bond cap stru wACC | 130% RESI ex di㈠ rences: CF135Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy ▼ ,_a. ars-函Merge & Center, $, % , 弼,8 C Paste B l u. Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard SB226 RZ Font Alignment Number Cells Edting SA SB SC CASE 3 SD SE SF SG SH SI SJ SK SM 226 227 228 229 230 231 232 233 234 235 236 237 238 239 240 241 242 243 244 1 KE CAPM , UTILITY, SH MILLER MODEL VU VALUE OF UNLEVERED FIRM EBIT(1-t)(1 -Ts)/ke(1-Ts) Ts = TAX ON STOCK INCOME-1096 Td = TAX ON DEBT INCOME-20% VU= 1000000*(1-0.4)(1-0, 10)/(0.15*(1-0, 10)) FOR UNLEVERED FIRM V - S-4000000 Ks = WACC = 15% VL = VU + DEBT[1-( 1-t)( 1-Ts)/( 1-Td)] 4000000 2000000(1 (1-0.4) (1-0.20)/(1-0.10)) 4933333.333 GAIN FROM LEVERAGE (MILLER MODEL) VL-VU- 4933333.33-4000000933333 GAIN FROM LEVERAGE (MM WITH TAX) VL-VU- 4800000-4000000- 800000 beta bond c DGM future INDEX INTL CAP BUDLEASING PV, FV, ANNUITYDIR cleanYIELD bond cap stru wACC | 130% RESI ex di㈠ rences: CF135